a. Perform binary addition on given numbers:  20 1/16 and 71/4 .  

b. Convert the following binary number into ASCII code and then into its equivalent decimal representation. 01001001010000010100110101011100 


1
Expert's answer
2021-05-25T13:05:32-0400

b.

binary to ASCII code:

"01001001010000010100110101011100=49\ 41\ 4D\ 5C" (HEX)

ASCII code: IAM\


binary to decimal:

"01001001010000010100110101011100=2^2+2^3+2^4+2^6+2^8+2^{10}+2^{11}+"

"+2^{14}+2^{16}+2^{22}+2^{24}+2^{27}+2^{30}=1,229,016,412"


a.

"20\ 1/16=20.0625=2^4+2^2+2^{-4}=1010.0001"

"71\/4=17.75=2^4+2^0+2^{-1}+2^{-2}=1001.11"

"1010.0001+1001.11=10011.1101"